OpenStudy (anonymous):

Choose the best answer choice. Answer A. 112.5 B. 67.5 C. 135 D. 56.25

OpenStudy (anonymous):



OpenStudy (ganpat):

i believe 85 + x = 180.. So, x = 180 - 85 = 95.. ??

OpenStudy (anonymous):